Use the following information for the remaining problems. First, construct an income statement and two balance sheets. Enter all answers as whole dollar numbers with no $ or commas (e.g. 20000). 2015 Sales = $60,000 2015 COGS = $25,000 2015 SG&A Expense = $10,000 2015 Depreciation Expense = $5,000 2015 Interest Expense = $3,000 Average Tax Rate = 30% Dividend Payout Ratio = 55% 2015 Current Assets = $24,000 and 2016 Current Assets = $27,000 2015 Net Working Capital = $5,000 Change in Net Working Capital = $1,000 2015 Total Fixed Assets = $100,000 2015 Accumulated Depreciation = $20,000 2015 Net Capital Spending = $12,000 2015 Long-term Debt = $40,000 2015 Common Stock = $22,000 2015 Cash Flow to Creditors = $1,000

What is 2016 Accumulated Retained Earnings?

What is 2015 Cash Flow to Shareholders?

What is 2016 Long-term Debt?

What is 2015 Free Cash Flow?
Answer #1

PROFIT AND LOSS ACCOUNT

particulars amount($) particulars amount($)
TO COGS 25000 BY SALES 60000
GROSS PROFIT (b/f) 35000
TOTAL 60000 TOTAL

60000

BY GROSS PROFIT(c/f) 35000
TO SG&A 10000
TO DEPRECIATION 5000
TO INTEREST EXPENSE 3000
NET PROFIT 17000
TOTAL 35000 TOTAL 35000

NET PROFIT =17000

AVERAGE TAX RATE =30%

PROFIT AFTER TAX=NET PROFIT -TAX AMOUNT

PAT=17000-(17000*30%)

PAT  =17000-5100=11900

BALANCE SHEET -2015

LIABILITIES AMOUNT($) ASSETS AMOUNT($)
COMMON STOCK 22000 FIXED ASSETS 107000
+NET PROFIT 17000 CURRENT ASSETS 24000
LONG TERM LIABILITIES 40000 CURRENT LIABILITIES 19000
TOTAL 150000 TOTAL 150000

NOTES : NET CAPITAL SPENDING=12000

=(ENDING VALUE OF FIXED ASSETS -BEGINNING VALUE OF FIXED ASSETS)+DEPRECIATION=12000

ENDING VALUE-100000+5000=12000

ENDING VALUE OF FIXED ASSETS=107000

1)ACCUMULATED RETAINED EARNINGS FOR YEAR 2016

NET PROFITS FOR YEAR 2015 =$17000

PROFIT AFTER TAX =$11900

DIVIDEND PAYOUT RATIO =55%

RETAINED EARNINGS =11900*55%=$6545

=(11900-6545)

=$5355

2) CASH FLOW TO SHARE HOLDERS FOR THE YEAR 2015

PROFIT AFTER TAXES = $ 11900

CASH FLOW = $11900*55%

=$6545

3)